How do I solve this?
0.1^n=100

Answer by richwmiller(9144)

You need to have noticed the powers of ten on each side.
.1=1/10
1/10=10(-1)
100=10^2
(10^(-1))^n=10^2
10^(-n)=10^2
n=-2